# DAISI: Data Assimilation with Inverse Sampling using Stochastic Interpolants

### v1

@@ D.4.4 Hyperparameters
High-dim. MMPS 100 0.03 0.4 1
@@ D.4.5 Results
Table 9: The RMSE for experiments on SQG and SEVIR. We display the mean and standard deviation across 10 independent trajectories, averaged over the last 20 (10 for SEVIR) steps. The best score for each experiment is highlighted in bold and the second best with an underline.
Experiment DAISI LETKF FlowDAS EnSF
Noisy $\mathbf{2.46}{\scriptstyle\pm 0.23}$ $\underline{2.51}{\scriptstyle\pm 0.26}$ $5.01{\scriptstyle\pm 0.69}$ $8.07{\scriptstyle\pm 1.01}$
Sparse $\mathbf{3.40}{\scriptstyle\pm 0.36}$ $\underline{4.55}{\scriptstyle\pm 0.53}$ $6.29{\scriptstyle\pm 0.90}$ $7.09{\scriptstyle\pm 0.74}$
Multimodal $\underline{3.71}{\scriptstyle\pm 0.95}$ $\mathbf{3.62}{\scriptstyle\pm 1.12}$ $6.93{\scriptstyle\pm 0.64}$ $9.64{\scriptstyle\pm 0.89}$
Saturating $\underline{2.96}{\scriptstyle\pm 0.17}$ $9.65{\scriptstyle\pm 0.69}$ $4.71{\scriptstyle\pm 0.69}$ $\mathbf{2.67}{\scriptstyle\pm 0.35}$
SEVIR $\mathbf{0.049}{\scriptstyle\pm 0.02}$ - $\underline{0.082}{\scriptstyle\pm 0.02}$ -
Table 10: The SSR for experiments on SQG and SEVIR. We display the mean and standard deviation across 10 independent trajectories, averaged over the last 20 (10 for SEVIR) steps. The best score for each experiment is highlighted in bold and the second best with an underline.
Experiment DAISI LETKF FlowDAS EnSF
Noisy $\textbf{1.06}{\scriptstyle\pm 0.03}$ $\underline{1.21}{\scriptstyle\pm 0.05}$ $0.78{\scriptstyle\pm 0.07}$ $0.49{\scriptstyle\pm 0.05}$
Sparse $1.24{\scriptstyle\pm 0.04}$ $\underline{1.23}{\scriptstyle\pm 0.10}$ $\textbf{1.13}{\scriptstyle\pm 0.06}$ $0.57{\scriptstyle\pm 0.05}$
Multimodal $1.66{\scriptstyle\pm 0.12}$ $\underline{0.76}{\scriptstyle\pm 0.15}$ $\textbf{1.21}{\scriptstyle\pm 0.07}$ $0.42{\scriptstyle\pm 0.04}$
Saturating $1.17{\scriptstyle\pm 0.03}$ $\underline{0.96}{\scriptstyle\pm 0.07}$ $\textbf{1.02}{\scriptstyle\pm 0.11}$ $1.60{\scriptstyle\pm 0.18}$
SEVIR $\underline{1.142}{\scriptstyle\pm 0.10}$ - $\textbf{1.085}{\scriptstyle\pm 0.13}$ -
@@ D.4.6 Ablation
One can replace the inversion step with a single SDEdit (meng2022sdedit) step, in which the forecast ${\bm{z}}_{1}$ is partially noised to the latent ${\bm{z}}_{t_{\min}}=\alpha_{t_{\min}}{\bm{z}}_{1}+\beta_{t_{\min}}{\bm{z}}_{0}$, ${\bm{z}}_{0}\sim\mathcal{N}(\bm{0},\mathbf{I})$. Using a tuned value of $t_{\min}=0.4$, this approach performs worse than using the backward SDE. The reason is that $t_{\min}$ simultaneously controls the strength of the corrective guidance and how much of the forecast information is preserved. This coupling introduces a trade-off, starting later preserves more forecast information but leaves less room for correction.
@@
Finally, we examine the effect of reduced assimilation frequency. Increasing the interval between observations amplifies forecast nonlinearities, which poses difficulties for LETKF. In the Noisy experiment, both LETKF and DAISI perform similarly with 3-hour assimilation intervals, but when the interval is extended to 12 hours, DAISI clearly outperforms LETKF (Figure 12).
@@ D.5 SEVIR
Here, we consider a real-life large-scale weather forecasting task using the Storm EVent Imagery and Radar (SEVIR) dataset (veillette2020sevir), which provides observations of severe convective storms across the continental United States. This specific experiment focuses on the Vertically Integrated Liquid (VIL) product, which serves as a 2-D proxy for precipitation intensity. Each data sample is a $128\times 128$ grid covering a $384\times$$384\text{\,}\mathrm{km}$ area at $2\text{\,}\mathrm{km}$ resolution, with snapshots recorded every $10\text{\,}\mathrm{min}$, for a total of $250\text{\,}\mathrm{min}$ (i.e., $25$ snapshots per sample).
For the forecast model, we used the checkpoint available in the official FlowDAS GitHub repository https://github.com/umjiayx/FlowDAS. This is based on a stochastic interpolant-based generative SDE for probabilistic forecasting, proposed in (chen2024probabilistic), with a U-Net backbone used for the drift. The model makes predictions autoregressively with inputs given by the past $6$ timesteps. For the guided FlowDAS baseline, we take the exact setup as in (chen2025flowdas), with parameters listed in Table 6. When beginning the assimilation, we use the first six ground truth states as initial conditions.
@@ Appendix E Additional figures
In Figures 1418 and 1924 we show additional samples and metrics for one of the assimilated trajectories.

### v2

@@ D.4.4 Hyperparameters
High-dim. MMPS 100 0.03 0.0 1
Non-stationary MMPS 100 0.03 0.5 3
@@ D.4.5 Results
Table 11: The RMSE for experiments on SQG and SEVIR. We display the mean and standard deviation across 10 independent trajectories, averaged over the last 20 (10 for SEVIR) steps. The best score for each experiment is highlighted in bold and the second best with an underline. Since SDA (smoothing) solves a different problem, we exclude it from the relative ranking.
Experiment DAISI LETKF FlowDAS EnSF SDA (filtering) SDA (smoothing)
Noisy $\underline{2.46}{\scriptstyle\pm 0.23}$ $2.51{\scriptstyle\pm 0.26}$ $5.01{\scriptstyle\pm 0.69}$ $8.07{\scriptstyle\pm 1.01}$ $\textbf{2.11}{\scriptstyle\pm 0.18}$ $2.15{\scriptstyle\pm 0.20}$
Sparse $\textbf{3.40}{\scriptstyle\pm 0.36}$ $4.55{\scriptstyle\pm 0.53}$ $6.29{\scriptstyle\pm 0.90}$ $7.09{\scriptstyle\pm 0.74}$ $\underline{4.17}{\scriptstyle\pm 0.44}$ $2.88{\scriptstyle\pm 0.20}$
Multimodal $\underline{3.71}{\scriptstyle\pm 0.95}$ $\textbf{3.62}{\scriptstyle\pm 1.12}$ $6.93{\scriptstyle\pm 0.64}$ $9.64{\scriptstyle\pm 0.89}$ $6.78{\scriptstyle\pm 0.70}$ $7.88{\scriptstyle\pm 0.52}$
Saturating $\underline{2.96}{\scriptstyle\pm 0.17}$ $9.65{\scriptstyle\pm 0.69}$ $4.71{\scriptstyle\pm 0.69}$ $\textbf{2.67}{\scriptstyle\pm 0.35}$ $7.52{\scriptstyle\pm 0.73}$ $6.73{\scriptstyle\pm 0.34}$
[…]
Table 12: The CRPS for experiments on SQG and SEVIR. We display the mean and standard deviation across 10 independent trajectories, averaged over the last 20 (10 for SEVIR) steps. The best score for each experiment is highlighted in bold and the second best with an underline. Since SDA (smoothing) solves a different problem, we exclude it from the relative ranking.
Experim. DAISI LETKF FlowDAS EnSF SDA (filtering) SDA (smoothing)
Noisy $\underline{1.32}{\scriptstyle\pm 0.13}$ $1.34{\scriptstyle\pm 0.14}$ $2.81{\scriptstyle\pm 0.43}$ $5.15{\scriptstyle\pm 0.78}$ $\textbf{1.13}{\scriptstyle\pm 0.10}$ $1.21{\scriptstyle\pm 0.11}$
Sparse $\textbf{1.73}{\scriptstyle\pm 0.18}$ $2.35{\scriptstyle\pm 0.28}$ $3.34{\scriptstyle\pm 0.53}$ $4.20{\scriptstyle\pm 0.56}$ $\underline{2.22}{\scriptstyle\pm 0.25}$ $1.53{\scriptstyle\pm 0.11}$
Multimodal $\textbf{1.81}{\scriptstyle\pm 0.44}$ $\underline{1.97}{\scriptstyle\pm 0.69}$ $3.66{\scriptstyle\pm 0.42}$ $6.38{\scriptstyle\pm 0.75}$ $3.88{\scriptstyle\pm 0.48}$ $3.82{\scriptstyle\pm 0.32}$
Saturating $\underline{1.54}{\scriptstyle\pm 0.09}$ $5.24{\scriptstyle\pm 0.35}$ $2.41{\scriptstyle\pm 0.39}$ $\textbf{1.33}{\scriptstyle\pm 0.16}$ $4.20{\scriptstyle\pm 0.44}$ $3.37{\scriptstyle\pm 0.14}$
[…]
Table 13: The SSR for experiments on SQG and SEVIR. We display the mean and standard deviation across 10 independent trajectories, averaged over the last 20 (10 for SEVIR) steps. The best score for each experiment is highlighted in bold and the second best with an underline. Since SDA (smoothing) solves a different problem, we exclude it from the relative ranking.
Experiment DAISI LETKF FlowDAS EnSF SDA (filtering) SDA (smoothing)
Noisy $\textbf{1.06}{\scriptstyle\pm 0.03}$ $\underline{1.21}{\scriptstyle\pm 0.05}$ $0.78{\scriptstyle\pm 0.07}$ $0.49{\scriptstyle\pm 0.05}$ $1.44{\scriptstyle\pm 0.04}$ $0.78{\scriptstyle\pm 0.02}$
[…]
Multimodal $1.66{\scriptstyle\pm 0.12}$ $0.76{\scriptstyle\pm 0.15}$ $\underline{1.21}{\scriptstyle\pm 0.07}$ $0.42{\scriptstyle\pm 0.04}$ $\textbf{0.91}{\scriptstyle\pm 0.07}$ $1.19{\scriptstyle\pm 0.04}$
Saturating $1.17{\scriptstyle\pm 0.03}$ $0.96{\scriptstyle\pm 0.07}$ $\textbf{1.02}{\scriptstyle\pm 0.11}$ $1.60{\scriptstyle\pm 0.18}$ $\underline{1.03}{\scriptstyle\pm 0.08}$ $4.81{\scriptstyle\pm 0.25}$
SEVIR ${1.19}{\scriptstyle\pm 0.12}$ $1.35{\scriptstyle\pm 0.14}$ ${1.09}{\scriptstyle\pm 0.17}$ $0.19{\scriptstyle\pm 0.08}$ $\textbf{1.03}{\scriptstyle\pm 0.10}$ $1.23{\scriptstyle\pm 0.09}$
@@ D.4.6 Ablation
One can replace the inversion step with a single SDEdit (Meng et al., 2022) step, in which the forecast ${\bm{z}}_{1}$ is partially noised to the latent ${\bm{z}}_{t_{\min}}=\alpha_{t_{\min}}{\bm{z}}_{1}+\beta_{t_{\min}}{\bm{z}}_{0}$, ${\bm{z}}_{0}\sim\mathcal{N}(\bm{0},\mathbf{I})$. Using a tuned value of $t_{\min}=0.4$, this approach performs worse than using the backward SDE. The reason is that $t_{\min}$ simultaneously controls the strength of the corrective guidance and how much of the forecast information is preserved. This coupling introduces a trade-off, starting later preserves more forecast information but leaves less room for correction.
@@
We also examine the effect of reduced assimilation frequency. Increasing the interval between observations amplifies forecast nonlinearities, which poses difficulties for LETKF. In the Noisy experiment, both LETKF and DAISI perform similarly with 3-hour assimilation intervals, but when the interval is extended to 12 hours, DAISI clearly outperforms LETKF (Figure 12).
Non-stationary observations:
Finally, we verify that DAISI can handle non-stationary observations, such as those arising from remote sensing instruments. The observation operator is linear with $\sigma_{\text{obs}}=1$ and consists of a band of width 4 px that shifts 4 px to the right at each timestep. We find that DAISI performs comparably to LETKF (Figure 13) while maintaining meaningful uncertainty (Figure 14).
@@
(b) The RMSE for each ensemble member at each time step.
(e) The spread of the ensemble at each time step.
(f) The spread skill ratio of the ensemble at each time step.
(g) The energy spectra at time step 1.
(h) The energy spectra at time step 50.
(i) The energy spectra at time step 100.
Figure 13: The results for the Non-stationary experiment.
Figure 14: A comparison of the ensemble mean, individual members, and ensemble standard deviation for each method at the last step of the assimilated trajectory for the Non-stationary experiment.
@@ D.5 SEVIR
Here, we consider a real-life large-scale weather forecasting task using the Storm EVent Imagery and Radar (SEVIR) dataset (Veillette et al., 2020), which provides observations of severe convective storms across the continental United States. This specific experiment focuses on the Vertically Integrated Liquid (VIL) product, which serves as a 2-D proxy for precipitation intensity. Each data sample is a $128\times 128$ grid covering a $384\times$$384\text{\,}\mathrm{km}$ area at $2\text{\,}\mathrm{km}$ resolution, with snapshots recorded every $10\text{\,}\mathrm{min}$, for a total of $250\text{\,}\mathrm{min}$ (i.e., $25$ snapshots per sample).
For the forecast model, we used the checkpoint available in the official FlowDAS GitHub repository https://github.com/umjiayx/FlowDAS. This is based on a stochastic interpolant-based generative SDE for probabilistic forecasting, proposed in (Chen et al., 2024), with a U-Net backbone used for the drift. The model makes predictions autoregressively with inputs given by the past $6$ timesteps. For the guided FlowDAS baseline, we take the exact setup as in (Chen et al., 2025), with parameters listed in Table 8. When beginning the assimilation, we use the first six ground truth states as initial conditions.
@@ Appendix E Additional figures
In Figures 1620 and 2126 we show additional samples and metrics for one of the assimilated trajectories.